OpenStudy (anonymous):

Find all solutions of the equation. (Enter each answer in the form θ + 2πk, 0 ≤ θ < 2π.) 1. sec(x) (2cos(x)+sqrt(3))=0 2. 2cos(2x)+1=0 If you can, please explain how you get the answer. Thanks!

zepdrix (zepdrix):

\[\large \sec x \left(2\cos x+\sqrt 3\right)=0\] For this first one let's recall that \(\large \sec x=\dfrac{1}{\cos x}\) If we distribute the secant x to each term in the brackets, what does it simplify down to?

OpenStudy (anonymous):

i got it right before u posted, thanks anyway

zepdrix (zepdrix):

oh ok